A smooth right circular cone of semi vertical angle alpha=tan^-1(5/12) is at rest on a horizontal plane .A rubber ring of mass 2.5kg which requires a force of 15N for an extension of 10cm is placed on the cone.Find the increase in the radius of the ring in equilibrium.

Grade:11

1 Answers

ksheer
5 years ago
  • Knsin(theta)=25
  • Kncos(theta)=15(x)
  • Tan(theta)=5/12
  • Increase in radius=10(x)/2
  • Believe me it's the right answer .
  • I am writing this because there should be more than 100 character s in my answer
  •